Convert the following hexadecimal IEEE-754 32-bit values to their decimal equivalents. Ox 40A80000 Oxc0000000
Please show all work
Answer #1

I hope it help.

thanks...

ANSWER:

Q 1).0x40A80000

Decimal from signed 2's complement:

1084751872

Binary number:

1000000101010000000000000000000

Decimal calculation:

(40A80000)₁₆ = (4 × 16⁷) + (0 × 16⁶) + (10 × 16⁵) + (8 × 16⁴) + (0 × 16³) + (0 × 16²) + (0 × 16¹) + (0 × 16⁰) = (1084751872)₁₀

Decimal number:

1084751872

Q 2). 0xC0000000

Decimal from signed 2's complement:

-1073741824

Binary number:

11000000000000000000000000000000

Decimal calculation:

(C0000000)₁₆ = (12 × 16⁷) + (0 × 16⁶) + (0 × 16⁵) + (0 × 16⁴) + (0 × 16³) + (0 × 16²) + (0 × 16¹) + (0 × 16⁰) = (3221225472)₁₀

Decimal number:

3221225472